OpenStudy (anonymous):

Evaluate the integral: (x^3+2-(5/x^2)) dx I need to see the steps to complete this problem, because I could not figure it out...

OpenStudy (anonymous):

\[\int\limits \left(x^3-\frac{5}{x^2}+2\right) \, dx\]

OpenStudy (anonymous):

Is that the integral?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You can break this integral into \[\int\limits x^3 \, dx-\int\limits \frac{5}{x^2} \, dx+\int\limits 2 \, dx\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you, I understand now

OpenStudy (anonymous):

Use this formula \[\int\limits x^N \, dx=\frac{1}{N+1}x^{N+1}+C\]

OpenStudy (anonymous):

(x^4/4)+(5/x)+(2x)+C

OpenStudy (anonymous):

Thank you!